OpenStudy (anonymous):

What is the solution of this linear system? 4x+y=-1 y=-2x+3

OpenStudy (anonymous):

well change the first to -4x-1=y -2x+3=y then you have to cancel out one of the variables so id multiply the top by -1 4x+1=-y -2x+3=y and solve like an addition equation 2x+4=0 2x=-4 x=-2 then solve for y

OpenStudy (anonymous):

Y is going to equal 7. or -7.

OpenStudy (anonymous):

it should just be 7

OpenStudy (anonymous):

-4(-2) +1= 7 -2(-2)+3=7

OpenStudy (anonymous):

What is the solution of this linear system? 2x + y = -4 y = -3x + 2

OpenStudy (radar):

Or you could rearrange the first equation to read; y = -4x - 1 We know that y = y so we can now write: -4x - 1 = -2x + 3 Now just solve for x -2x = 4 x = -2 you can solve for y by substituting that value for x in either of the two original equations.

OpenStudy (anonymous):

well you would do the same steps so y= -3x+2 y= -2x-4 turn one negative to cancel out y= -3x+2 -y= 2x+4 x+6=0 x=-6 -3(-6)+2= 20 y=20

OpenStudy (anonymous):

Thanks everyone!